# Readers Prefer Outputs of AI Trained on Copyrighted Books over Expert Human Writers

### v2

@@ 18 Ian McEwan (✓) 36 Percival (✓)Everett
As can be seen in Figure11 writing quality has a strong correlation with a faithful style/voice fidelity for In-context Prompting condition compared to Finetuning. The high expert correlation for In-context Prompting suggests that models cannot emulate an author’s style properly just by prompting and that experts use stylistic cues as a major factor in quality judgments - likely detecting “AI-ness” (clichés, purple prose, too much exposition, lack of subtext, mixed metaphors) in the style that makes them rate quality lower too. With finetuning, experts can appreciate quality independent of stylistic fidelity, suggesting the as AI loses its tell-tale signs experts can evaluate each dimension on its own merits. For lay readers the correlation only drops slightly from 0.52 (In-context prompting) to 0.45 (fine-tuned), compared to experts’ dramatic drop from 0.64 to 0.39. This also mean that lay readers might focus on more surface-level features such as grammar, clarity or flow rather than subtle stylistic markers that distinguishes AI from human writing.
@@
Figure 12,13, 14 and 15 shows emulations of the original excerpt written by Colson Whitehead, Ottessa Moshfegh, Jonathan Franzen and Han Kang. There are 3 distinct MFA emulations per author (MFA1, MFA2, MFA3). For the In-context Prompting set up we prompt Claude3.5 Sonnet (AI1), Gemini1.5 (AI2) and GPT-4o(AI3) to produce excerpts that can be pitted against MFA written excerpts. In Finetuning setup we finetune GPT4-o and compare the fine-tuned output AI4 against MFA1, MFA2, MFA3. The output from fine-tuned GPT4-o (AI4) was generated using default hyperparameters.
@@
Figure 17 and 18 shows the writing quality preference evaluation results from expert readers. Expert preferences are often grounded in solid reasoning and they often agree on similar snippets to support their reasoning. For instance look at uncooked spaghetti as an overwrought/awkward metaphor highlighted by both Expert Reader 1 and 3 while tarot card predicting his fate highlighted by Expert Reader 2 and 3. Grounding preferences in reasoning helps us uncover their mental models. It is also worth noticing that the preference often shifts towards AI once it is fine-tuned on authors entire oeuvre. Fine-tuning helps get rid of the officious disembodied robovoice that is characteristic of ChatGPT. Here experts praise the model for idiosyncratic narrative voice. By default due to post training guardrails GPT4-o generates a rather safe and somewhat polite text in the in-context prompting set up. But what is particularly impressive here is when fine-tuned the vocabulary of the model generated text veers towards ribald and somewhat profane text that is characteristic of Tulathimutte’s voice (particularly Rejection from where the text is drawn). In Figure 19 we see that experts prefer the MFA emulation as more faithful to Junot Diaz’s voice. In fact the excerpt written by Gemini-1.5-Pro is oddly poetic and rife with purple prose and cliches which is nothing like Junot Diaz’s voice that is often marked by a dazzling hash of Spanish, English, slang, literary flourishes, and pure virginal dorkiness as noted by all three expert readers. However when fine-tuned on authors’ entire oeuvre (Figure 20) we see that the model learns these references and uses them in a rather wry and humorous manner (papi chulo, fifty pendejas, little pito off, las plagas) that is appreciated by expert readers leading them to prefer AI over MFAs.
Figure 17: MFA vs AI emulation of Tony Tulathimutte, readerd by 3 expert readers all of whom agree that MFA written excerpt is superior in terms of writing quality.
Figure 18: MFA vs AI emulation of Tony Tulathimutte, judged by 3 expert readers all of whom agree that Fine-tuned AI written excerpt is superior in terms of writing quality
@@
Figure 20: MFA vs AI emulation of Junot Diaz, judged by 3 expert readers all of whom agree that Fine-tuned AI written excerpt is superior in terms of stylistic fidelity
@@
Figure 21: Pangram AI likelihood scores vs GPTZero AI Likelihood scores for MFA, In-context and Fine-tuned
@@
style In_Context GPT4o_baseline Expert 150
style In_Context GPT4o_baseline Lay 250
@@
Table 4: Observation counts for stylistic fidelity in the fine-tuned setting. Balanced counts (Human = GPT4o_fine-tuned) reflect the one-to-one comparison design for H2.
@@
style Fine_tuned GPT4o_fine-tuned Expert 270
style Fine_tuned GPT4o_fine-tuned Lay 450
@@
quality In_Context GPT4o_baseline Expert 150
quality In_Context GPT4o_baseline Lay 250
@@
quality Fine_tuned GPT4o_fine-tuned Expert 270
quality Fine_tuned GPT4o_fine-tuned Lay 450
@@
GPT4o_baseline GPT-4o in-context prompting AI excerpt; excerpt_model $\in$ {GPT4o}; setting = In_Context
@@
GPT4o_fine-tuned GPT-4o fine-tuned AI excerpt; excerpt_model = GPT4o_Fine-tuned
@@
writer_typeGPT4o_baseline -1.655 0.402 -4.120 3.78e-05
@@
writer_typeGPT4o_baseline:reader_typeLay 1.468 0.446 3.292 9.95e-04
@@
writer_typeGPT4o_fine-tuned 2.100 0.282 7.435 1.04e-13
@@
writer_typeGPT4o_fine-tuned:reader_typeLay 0.015 0.372 0.042 0.967
@@
writer_typeGPT4o_baseline -2.406 0.473 -5.083 3.71e-07
@@
writer_typeGPT4o_baseline:reader_typeLay 3.065 0.518 5.919 3.23e-09
@@
writer_typeGPT4o_fine-tuned 0.627 0.244 2.571 0.0102
@@
writer_typeGPT4o_fine-tuned:reader_typeLay 0.258 0.314 0.821 0.412
@@
style In_Context GPT4o_baseline:reader_typeLay 1.468 0.446 3.292 9.95e-04
@@
style Fine_tuned GPT4o_fine-tuned:reader_typeLay 0.015 0.372 0.042 0.967
quality In_Context GPT4o_baseline:reader_typeLay 3.065 0.518 5.919 3.23e-09
@@
quality Fine_tuned GPT4o_fine-tuned:reader_typeLay 0.258 0.314 0.821 0.412
@@
All code and data necessary to reproduce Figures 2-4 and the associated statistical analyses will be made publicly available on paper acceptance. Please contact authors if you need it:

### v3

@@ 18 Ian McEwan (✓) 36 Percival (✓)Everett
As can be seen in 14 writing quality has a strong correlation with a faithful style/voice fidelity for In-context Prompting condition compared to Finetuning. The high expert correlation for In-context Prompting suggests that models cannot emulate an authors style properly just by prompting and that experts use stylistic cues as a major factor in quality judgments - likely detecting ”AI-ness” (clichés, purple prose, too much exposition, lack of subtext, mixed metaphors) in the style that makes them rate quality lower too. With finetuning, experts can appreciate quality independent of stylistic fidelity, suggesting the as AI loses its tell-tale signs experts can evaluate each dimension on its own merits. For lay readers the correlation only drops slightly from 0.52 (In-context prompting) to 0.45 (finetuned), compared to experts’ dramatic drop from 0.64 to 0.39. This also mean that lay readers might focus on more surface-level features such as grammar, clarity or flow rather than subtle stylistic markers that distinguishes AI from human writing.
@@
Figure 15,16, 17 and 18 shows emulations of the original excerpt written by Colson Whitehead, Ottessa Moshfegh, Jonathan Franzen and Han Kang. There are 3 distinct MFA emulations per author (MFA1, MFA2, MFA3). For the In-context Prompting setup we prompt Claude 3.5 Sonnet (AI1), Gemini 1.5 Pro (AI2) and GPT-4o(AI3) to produce excerpts that can be pitted against MFA written excerpts. In finetuning setup we finetune GPT-4o and compare the fine-tuned output AI4 against MFA1, MFA2, MFA3. The output from fine-tuned GPT-4o (AI4) was generated using default hyperparameters.
@@
Figure 20 and 21 shows the writing quality preference evaluation results from expert readers. Expert preferences are often grounded in solid reasoning and they often agree on similar snippets to support their reasoning. For instance look at uncooked spaghetti as an overwrought/awkward metaphor highlighted by both Expert Reader 1 and 3 while tarot card predicting his fate highlighted by Expert Reader 2 and 3. Grounding preferences in reasoning helps us uncover their mental models. It is also worth noticing that the preference often shifts towards AI once it is finetuned on authors entire oeuvre. Fine-tuning helps get rid of the officious disembodied robovoice that is characteristic of ChatGPT. Here experts praise the model for idiosyncratic narrative voice. By default due to post-training guardrails GPT-4o generates a rather safe and somewhat polite text in the in-context prompting setup. But what is particularly impressive here is when finetuned the vocabulary of the model generated text veers towards ribald and somewhat profane text that is characteristic of Tulathimutte’s voice (particularly Rejection from where the text is drawn). In Figure 22 we see that experts prefer the MFA emulation as more faithful to Junot Diaz’s voice. In fact the excerpt written by Gemini-1.5-Pro is oddly poetic and rife with purple prose and cliches which is nothing like Junot Diaz’s voice that is often marked by a dazzling hash of Spanish, English, slang, literary flourishes, and pure virginal dorkiness as noted by all three expert readers. However when finetuned on authors’ entire oeuvre (Figure 23) we see that the model learns these references and uses them in a rather wry and humorous manner (papi chulo, fifty pendejas, little pito off, las plagas) that is appreciated by expert readers leading them to prefer AI over MFAs.
Figure 20: MFA vs AI emulation of Tony Tulathimutte, read by 3 expert readers all of whom agree that MFA written excerpt is superior in terms of writing quality.
Figure 21: MFA vs AI emulation of Tony Tulathimutte, judged by 3 expert readers all of whom agree that Finetuned AI written excerpt is superior in terms of writing quality
@@
Figure 23: MFA vs AI emulation of Junot Diaz, judged by 3 expert readers all of whom agree that Finetuned AI written excerpt is superior in terms of stylistic fidelity
@@
Figure 24: Pangram AI likelihood scores vs GPTZero AI Likelihood scores for MFA, In-context and Finetuned
@@
style In_Context GPT-4o_baseline Expert 150
style In_Context GPT-4o_baseline Lay 250
@@
Table 4: Observation counts for stylistic fidelity in the fine-tuned setting. Balanced counts (Human = GPT-4o_finetuned) reflect the one-to-one comparison design for H2.
@@
style Fine_tuned GPT-4o_finetuned Expert 270
style Fine_tuned GPT-4o_finetuned Lay 450
@@
quality In_Context GPT-4o_baseline Expert 150
quality In_Context GPT-4o_baseline Lay 250
@@
quality Fine_tuned GPT-4o_finetuned Expert 270
quality Fine_tuned GPT-4o_finetuned Lay 450
@@
GPT-4o_baseline GPT-4o in-context prompting AI excerpt; excerpt_model $\in$ {GPT-4o}; setting = In_Context
@@
GPT-4o_finetuned GPT-4o fine-tuned AI excerpt; excerpt_model = GPT-4o_Finetuned
@@
writer_typeGPT-4o_baseline -1.655 0.402 -4.120 3.78e-05
@@
writer_typeGPT-4o_baseline:reader_typeLay 1.468 0.446 3.292 9.95e-04
@@
writer_typeGPT-4o_finetuned 2.100 0.282 7.435 1.04e-13
@@
writer_typeGPT-4o_finetuned:reader_typeLay 0.015 0.372 0.042 0.967
@@
writer_typeGPT-4o_baseline -2.406 0.473 -5.083 3.71e-07
@@
writer_typeGPT-4o_baseline:reader_typeLay 3.065 0.518 5.919 3.23e-09
@@
writer_typeGPT-4o_finetuned 0.627 0.244 2.571 0.0102
@@
writer_typeGPT-4o_finetuned:reader_typeLay 0.258 0.314 0.821 0.412
@@
style In_Context GPT-4o_baseline:reader_typeLay 1.468 0.446 3.292 9.95e-04
@@
style Fine_tuned GPT-4o_finetuned:reader_typeLay 0.015 0.372 0.042 0.967
quality In_Context GPT-4o_baseline:reader_typeLay 3.065 0.518 5.919 3.23e-09
@@
quality Fine_tuned GPT-4o_finetuned:reader_typeLay 0.258 0.314 0.821 0.412
@@
All code and data necessary to reproduce Figures 2-4 and the associated statistical analyses are publicly available: